### v1

# From negative to positive $\Lambda$ through cosmological decreasing temperatures and its connection to spacetime foliation and string theory

Abstract
String theories naturally predict a negative, while observations on the exponential expansion of the present Universe requires a positive value for the cosmological constant $\Lambda$. Solution to resolve this discrepancy is known in the framework of string theory and given for Asymptotically Safe (AS) quantum gravity by using Renormalization Group (RG). However, the former might describes unstable worlds and the latter lacks the inclusion of temperature which is otherwise crucial in the early Universe. Here we present a proposal for resolving this conflict by using a modified thermal RG method where the temperature parameter $T$ is given by the inverse radius of the compactified time-like dimension, similarly to spacetime foliation. In our scenario not the dimensionful temperature $T$, but the dimensionless temperature $\tau=T/k$ is kept constant when the RG scale $k$ is sent to zero and string theory is assumed to take place at very high while AS quantum gravity at intermediate and low temperatures. We show that the modified thermal RG study of AS quantum gravity models at very high temperatures results in a negative cosmological constant while turns it into a positive parameter for low temperatures.
1 Introduction
In this work we present a proposal for resolving the conflict between string theories and late time cosmology regarding the sign of the cosmological constant. Indeed, the negative value of the cosmological constant was naturally predicted by string theories, which contradicted the need for a positive value based on current cosmological observations. This motivated attempts to resolve the discrepancy in the framework of string theory [1] by wrapping antibranes, but to get the required small and positive cosmological constant one has to wrap many fluxes and it turned out that these solutions might describe unstable worlds [2].
Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a functional Renormalization Group (RG) method [3]. The concept was to assume string theory at the ultraviolet (UV) and Asymptotically Safe (AS) quantum gravity & matter at the infrared (IR) scaling regimes connected by the RG scale $k$ which serves as a bridge between the negative UV and the positive IR values of the cosmological constant. However, the RG scale $k$ is introduced artificially in the Wilsonian approach [4] and the quantized theory must be obtained in the physical limit $k\to 0$, so its use is not fully justified to connect early and late time cosmologies. In addition, temperature, which is a relevant parameter in cosmology, is missing in this approach.
Thus, here we suggest to use the temperature instead of the RG scale $k$ to connect string theory at very high temperature and AS quantum gravity at intermediate and low temperatures of the expanding Universe. To achieve this one has to extend the zero-temperature functional RG method [5] to finite temperature. This can be done by using the standard finite temperature extension of QFT where the time-like dimension is compactified [6, 7, 8] and its inverse radius plays the role of the dimensionful temperature parameter $T$ (in natural units). However, the usual choice of the perturbative RG approach, when the temperature is linked to the running (perturbative) RG scale $\mu$, i.e., $T=\mu/(2\pi)$, cannot be used in the functional RG method because one has to take the limit $k\to 0$ to obtain the quantized theory. Therefore, in [9] the temperature $T$ is kept constant over the RG flow, i.e., it is linked to a fixed (not running) momentum scale. Most of the further literature on thermal functional RG is based on this assumption. In this fixed $T$ approach, the dimensionful temperature is well defined and the limit $k\to 0$ can be done safely, but there is a price to pay: the RG flow equations have no fixed point solutions. Therefore, progress in this direction has been hampered by the fact that the usual extension of thermal functional RG method at finite temperature does not reproduce non-trivial fixed points needed for the RG analysis of various questions formulated in the framework of thermal QFT.
To overcome these difficulties, in Refs. [10, 11, 12] we proposed a modification of the usual functional RG approach af finite temperature by relating the temperature parameter to the running RG scale as $T\equiv k_{T}=\tau k$ (in natural units). In our approach
$\tau=T/k$
is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale, so it is not considered as the temperature, but rather understood as a running cutoff for thermal fluctuation. The dimensionless $\tau$ is identified as the physical temperature of the surrounding plasma. In this case the RG flow equations have real (and not pseudo) fixed points, which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations. In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points. In our fixed $\tau$ approach the dimensionful temperature is not defined, but the fixed points are. This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [12] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme. Moreover, in the recent work [11] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ aproach. The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [13] (i.e., non-Gaussian UV) fixed point. We showed that in the high temperature limit ($\tau\to\infty$) the $g$-coordinate of the Reuter fixed point vanishes and the cosmological constant takes on a negative value in the limit $k\to 0$. It is not in disagreement with observations, since during the thermal evolution of the Universe a thermal phase transition occurs and the cosmological constant runs to the expected positive value at low temperatures. This mechanism to solve the sign-problem of the cosmological constant is represented on Fig. 1.
Figure 1: Schematic figure of the time evolution (temperature, i.e., $\tau$-dependence) of the scale factor $R(t)$ and the cosmological constant $\Lambda(t)$.
Motivated by the results in [11], here we study the generality of these findings. We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [14, 15]) has the cosmological constant with a negative (positive, repsectively) sign for large (small) temperature. Therefore, the scheme represented in Fig. 1 can be used in every case which can be seen as a natural solution for the sign problem of the cosmological constant. In particular, we study three extensions/variants of the simplest AS quantum gravity: the conformally reduced and the ghost-improved versions and its extension by scalar matter fields. In addition to that we discuss the connection to spacetime foliation.
2 CREH gravity at $T=0$
As a first step, we summarise the main ideas behind the Conformally Reduced Einstein-Hilbert (CREH) truncated gravity at zero temperature [16]. In this model the approximation of Quantum Einstein Gravity (QEG) RG flow is done in two steps. Firstly, one takes the usual EH truncation, then one performs the conformal reduction, where only the conformal factor of the metric is quantized. The parametrization of the conformal factor is done in terms of the $\phi(x)$ scalar function (with kinetic term $\sim(\partial_{\mu}\phi)^{2}$). In $d=4$ dimensions the metric is given as $g_{\mu\nu}=\phi^{2}\widehat{g}_{\mu\nu}$, where $\widehat{g}_{\mu\nu}$ is the non-dynamical reference metric. The EH action reads
$S_{\rm EH}[g_{\mu\nu}]=\frac{1}{16\pi G}\int d^{4}x\,\sqrt{g}\,(2\Lambda-\mathcal{R}(g)),$ (1)
where $G$ is the Newton constant, $\Lambda$ is the cosmological constant, $\mathcal{R}$ is the Ricci scalar and $g=\text{det}(g_{\mu\nu})$. Performing Weyl rescalings leads to a $\phi^{4}$-like theory with $S_{\rm EH}=\int d^{4}x\,\sqrt{\widehat{g}}{\cal L}_{\rm EH}$, with
${\cal L}_{\rm EH}=-\frac{3}{4\pi G}\,\bigg(\frac{1}{2}\widehat{g}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\,+\frac{1}{12}\widehat{\mathcal{R}}\phi^{2}-\frac{1}{6}\Lambda\phi^{4}\bigg).$ (2)
Since the kinetic term is negative in Eq. (2), a rapidly varying $\phi(x)$ could cause $S_{\rm EH}$ to become arbitrary negative, this being called the conformal factor instability. One can introduce an inverted action ($S_{\rm inv}\equiv-S_{\rm EH}$) to shift the negative sign to the potential term which is used in the path integral. By using the RG formalism to the background field approach, see  A, leads then to the effective action $\Gamma_{k}[\bar{f};\chi_{B}]=\int d^{4}x\,\sqrt{\widehat{g}}{\cal L}^{\rm CREH}_{k}$ with
${\cal L}^{\rm CREH}_{k}=-\frac{3}{4\pi G_{k}}\left(-\frac{1}{2}\phi\widehat{\Box}\phi+\frac{1}{12}\widehat{\mathcal{R}}\phi^{2}-\frac{1}{6}\Lambda_{k}\phi^{4}\right),$ (3)
where $\widehat{\Box}=\widehat{g}^{-1/2}\partial_{\mu}\widehat{g}^{1/2}\widehat{g}^{\mu\nu}\partial_{\nu}$ denotes the Laplace-Bertrami operator belonging to the reference metric. In the RG flow equations dimensionless couplings are used, defined as
$g_{k}=k^{2}G_{k},\hskip 28.45274pt\lambda_{k}=k^{-2}\Lambda_{k}\,.$ (4)
leading to the following beta-functions
$\beta_{g}=$ [ $\,[\,2+\eta_{N}]\,g_{k}\,$, (5)
$\beta_{\lambda}=$ $\,-(2-\eta_{N})\,\lambda_{k}+\frac{g_{k}}{2\pi}\,[\,\Phi_{2}^{1}(-2\lambda_{k})-\frac{1}{2}\,\eta_{N}\,\tilde{\Phi}_{2}^{1}(-2\lambda_{k})\,],$
with threshold functions $\Phi_{2}^{1}(w)$, $\tilde{\Phi}_{2}^{1}(w)$ and anomalous dimension $\eta_{N}$, see  A.
3 QEG with matter at $T=0$
We summarise briefly the RG study of QEG coupled to $N$-component scalar field at zero temperature [17]. The model is interesting for the influence of the number of scalars and the scalar gravitational coupling on the flow of the Newtonian and the cosmological constant. One has ${\cal S}_{\rm N}=\int d^{4}x\sqrt{-g}{\cal L}_{\rm N}$, with the Lagrangian density
${\cal L}_{\rm N}=\Bigg[\frac{(-\mathcal{R}+2\Lambda)}{16\pi G}+\frac{1}{2}\partial_{\mu}\phi^{i}\partial^{\mu}\phi_{i}+\frac{1}{2}\xi\mathcal{R}\phi^{i}\phi_{i}\Bigg],$ (6)
where i=1,,Ni=1,\,...,\,N and $\xi$ is the scalar gravitational coupling. The background field method is used during the derivation of the beta-functions, see  A, resulting in
$\beta_{g}=$ [ $\,[\,2+\eta_{N}(k)\,]\,g_{k}\,$, (7)
$\beta_{\lambda}=$ [ $-[\,2-\eta_{N}(k)\,]\lambda_{k}+\frac{1}{2\pi}g_{k}\,[\,10\,\Phi_{2}^{1}(-2\lambda_{k})\,+$
$+\,(N-8)\,\Phi_{2}^{1}(0)-5\,\eta_{N}(k)\,\tilde{\Phi}_{2}^{1}(-2\lambda_{k})\,].$
4 Ghost-improved EH gravity at $T=0$
Lastly, let us study the EH truncated gravity at zero temperature with quantum effects captured by the wave-function renormalization $Z_{k}^{c}$, which multiplies the ghost-kinetic term. One obtains the EH truncated gravity [18] by fixing $Z_{k}^{c}=1$. We follow the analysis of the ghost-improved model of [19], where the motivation stems from a QCD analogy, $Z_{k}^{c}$ playing an important role in the IR theory [20, 21, 22, 23], and the computation revealing the interplay between gravitational beta-functions and ghosts.
The ansatz for the scale dependent effective action is
$\Gamma_{k}[g,C,\bar{C};\bar{g},c,\bar{c}]=$
$\Gamma_{k}^{\text{grav}}[g]\,+\,\Gamma_{k}^{\text{gf}}[g;\bar{g}]\,+\,\Gamma_{k}^{\text{gh}}[g,C,\bar{C};\bar{g},c,\bar{c}]\,$, (8)
where $C,\bar{C}$ are the classical ghost fields, and $c,\bar{c}$ are their associated background fields: one has $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, $\bar{C}_{\mu}=\bar{c}_{\mu}+\bar{f}_{\mu}$ and $C_{\mu}=c_{\mu}+f_{\mu}$, where $h_{\mu\nu}$, $f_{\mu}$, $\bar{f}_{\mu}$ mark the expectation value of the quantum fluctuations around the background. The ansatz is constructed from the $\Gamma_{k}^{\text{grav}}$ gravitational term, the $\Gamma_{k}^{\text{gf}}$ gauge-fixing term and the $\Gamma_{k}^{\text{gh}}$ ghost term, with their explicit forms also given in [19]. The harmonic gauge choice is also used in order to compare the result to the EH truncated gravity without ghost-improvement. Beta-functions in $d$ dimensions are found to be
$\beta_{g}=$ $\,(d-2+\eta_{N})\,g_{k}\,$, (9)
$\beta_{\lambda}=$ $-(2-\eta_{N})\lambda_{k}+\frac{1}{2}g_{k}(4\pi)^{1-d/2}$
$\Big[2d(d+1)\Phi^{1,0}_{d/2}(-2\lambda_{k})-8d\Phi^{1,0}_{d/2}(0)$
$-d(d+1)\eta_{N}\tilde{\Phi}^{1,0}_{d/2}(-2\lambda_{k})+4d\eta_{c}\tilde{\Phi}^{1,0}_{d/2}(0)\Big],$
with threshold functions $\Phi_{n}^{p,q}(w)$, $\tilde{\Phi}_{n}^{p,q}(w)$,  A.
The wave-function renormalization of ghosts $Z_{k}^{c}$ is completely determined by $g_{k}$ and $\lambda_{k}$, since it enters the beta-function through the ghost anomalous dimension $\eta_{c}=-\partial_{t}\,\text{ln}\,Z_{k}^{c}$.
5 AS gravity at finite temperature
In order to generalize the $T=0$ gravitational models discussed in the previous sections to their finite temperature counterpart, one has to extend the zero temperature RG method to finite temperature. The formulation is done in Euclidean spacetime, which can be achieved by Wick rotation, i.e., $t\to-it_{E}$, where $t_{E}$ is the Euclidean time. This transforms the action as $\int_{0}^{t}dt\int d\,^{d-1}x\,\mathcal{L}\to\int_{0}^{\beta}dt_{E}\int d\,^{d-1}x\,\mathcal{L}(t\to-it_{E})$. Bosonic fields obey periodic boundary conditions, with $\beta=it=1/T$ periodicity. Finite temperature QFT requires the modification of the momentum integral as
$\int\frac{d\,^{d}p}{(2\pi)^{d}}\hskip 5.69054pt\to\hskip 5.69054ptT\sum_{m}\int\frac{d\,^{d-1}p}{(2\pi)^{d-1}}$ (10)
with the Matsubara summation over $m\in\mathbb{Z}$ exchanging one of the momentum integrals. This implies that one has to replace the zeroth component of the momentum with $\omega_{m}$ Matsubara frequencies, i.e., $p^{2}\to p^{2}+\omega_{m}^{2}$, where $\omega_{m}=2\pi mT$ holds for bosonic frequencies.
Due to the periodic boundary condition, the finite temperature QFT is described on a cylindrical spacetime, with radius $R=1/T$. For the reasons discussed in detail in the Introduction, we chose to link $T$ to the scale $k$ as $T=\tau k$.
Introducing the Matsubara summation (10) and implementing our temperature relation $T=\tau k$ implies the following changes to the the zero-temperature threshold functions (see  A for details):
$\Phi^{p}_{n}(w,\tau)=\frac{\,2\tau\sqrt{\pi}}{\Gamma\left(n-\frac{1}{2}\right)}\ \sum_{m\ =\ -\infty}^{\infty}\int^{\infty}_{0}dy\ y^{n-\frac{3}{2}}$
$\frac{R(y)-y\,R^{\prime}(y)}{\left[y+(2m\pi\tau)^{2}+R(y)+w\right]^{p}},$ (11)
$\tilde{\Phi}^{p}_{n}(w,\tau)=\frac{\,2\tau\sqrt{\pi}}{\Gamma\left(n-\frac{1}{2}\right)}\ \sum_{m\ =\ -\infty}^{\infty}\int^{\infty}_{0}dy\ y^{n-\frac{3}{2}}$
$\frac{R(y)}{\left[y+(2m\pi\tau)^{2}+R(y)+w\right]^{p}}.$ (12)
Subsequently, these expressions are introduced into the beta-functions of the gravitational models, the only exception being the ghost-improved scenario, in which the generalized threshold functions are used which requires an additional $y+R(y)]^{q}\to[y+(2m\pi\tau)^{2}+R(y)]^{q}$ replacement.
In all three gravitational models we found that the $g$-component of the Reuter fixed point ($g^{*}$) disappears as the temperature increases, i.e., with $\tau\to\infty$, see Fig. 2.
Figure 2: Thermal RG flow diagrams of CREH, QEG with $N$-component scalar field (with chosen parameters $\xi=10$ and $N=2$) and Ghost-improved EH gravity for various $\tau$ dimensionless temperature values. With $\tau\to\infty$ the $g^{*}\to 0$ limit is reached in each case.
Additionally, the slope of the separatrix decreases with increasing $\tau$ in each case and, in the limit $\tau\to\infty$, only the $\Lambda<0$ phase survives. This result shows that AS quantum gravity naturally predicts a negative cosmological constant for high temperatures. In the early (high-temperature) Universe – whether one considers CREH, QEG with $N$-component scalar field, or ghost-improved gravity – starting from particular initial condition in the $\Lambda<0$ phase [11], the decreasing temperature of the Universe drives the system into the $\Lambda>0$ phase, which is the one consistent with current observations, see Fig. 3.
The system can undergo either a QPT at a fixed finite temperature or a CPT can occur at a fixed quantum parameter. There is some freedom in the choice of the parameters, since at $\tau\to\infty$ the $\lambda$-component of the Reuter fixed point ($\lambda^{*}$) is essentially constant in all of the discussed models, as seen in the inset of Fig. 3. This implies that the slope of the separatrix $^{*}/\lambda^{*}, gg^{*}$ itself, or $g^{*}\lambda^{*}$ are all adequate candidates for quantum parameters. We chose the latter option, for it is a dimensionless combination of the couplings in $d=4$, i.e., $G_{k}\Lambda_{k}=g_{k}\lambda_{k}$ applies. Since $g^{*}\to 0$ is only achieved in the $\tau\to\infty$ limit, it is worth to investigate their combination as the quantum parameter vanishes. As it turns out, $\tau g^{*}$ reaches a constant value in the $g^{*}\lambda^{*}\to 0$ limit, which provides the required form for the QPT-CPT diagram within all three models, as seen in Fig. 3.
Figure 3: QPT-CPT diagram of various AS gravity models in terms of the dimensionless temperature $\tau$ and the $g$-coordinate of the Reuter fixed point. Black lines, i.e., the function $g^{\star}(\tau_{c})$, are critical lines which separate the $\lambda<0$ and the $\lambda>0$ phases. For a given (but fixed) $g^{\star}$-value, for $\tau>\tau_{c}$ or $\tau<\tau_{c}$ the particular model is in the $\lambda<0$ or its $\lambda>0$ phase. The inset shows how the positions of the Reuter fixed point ($g^{\star}$, $\lambda^{\star}$) changes by $\tau$ in case of the three variants.
In all three models one finds non-computable regions with boundaries signaled by the divergence of the anomalous dimension which expands with increasing $\tau$. In CREH gravity, at zero temperature this pole appears at $\lambda=0.5$, however, in the limit $\tau\to\infty$ it is shifted to $\lambda\simeq 0.3$ for non-zero $g$ but $\lambda=0.5$ for vanishing $g$.
6 Connection to spacetime foliation
Our modified thermal RG framework shows many formal similarities to RG methods implementing foliated spacetimes within the Arnowitt-Deser-Misner (ADM) formalism [24, 25], as summarised in  B, based on Refs. [26, 27, 28, 29, 30].
The formulation for the RG equation, where the gravitational degrees of freedom are carried by the ADM fields has been constructed in Refs. [28, 29]. In these works the RG flow captured by the ADM-decomposed EH action was studied, revealing that the beta-functions parametrically depend on the dimensionless Matsubara (or Kaluza-Klein) mass $m$, which is related to the size of the time direction $R$ and it is defined as $m=2\pi/(Rk)$. If we implement the Matsubara formalism $m=2\pi T/k$ applies, hence the dimensionful temperature takes the form $T=mk/(2\pi)$. Comparing this to our key identification $T=\tau k$, one can see that the dimensionless temperature $\tau$ serves the same purpose as the Matsubara mass $m$. Figure 2 of Ref. [29] is also consistent with this observation, as the position of the Reuter fixed point exhibits a strikingly similar dependence on $m$ to that shown in the inset of Fig. 3. The beta-function for $m$ was studied in the work mentioned previously, and the $\beta_{m}(g,\lambda,m)=\partial_{t}m_{k}=0$ approximation was taken. Additionally, in [31] both constant and running $m$ were studied in relation to Hořava–Lifshitz gravity. This approach might be appropriate regarding foliated spacetimes, but in case of finite temperature formalism we advocate for taking $m$, i.e., $\tau$, as constant during the RG transformations. In this way, temperature fluctuations are integrated out in a manner similar to quantum fluctuations which makes possible to draw the QPT-CPT diagrams of QFT models.
7 Conclusions
The main result of this work is a prediction: if one assumes that AS gravity is a viable theoretical framework to connect early and late time cosmologies, then based on our thermal RG analysis, the cosmological constant must be negative in the early and positive in the present Universe. The latter is in agreement with observations on the accelerated expansion of the Universe at present while the former needs a verification or falsification. However, independently of possible future tests on the negative cosmological constant of the early Universe, our result receives an important application in string theories which naturally produce the cosmological constant with a negative value in agreement with the prediction of this work.
To support the above general statement, in this work we have performed the thermal RG study of three variants of the simplest (EH truncated) AS quantum gravity. In all these cases we have found the same picture: a negative value for the cosmological constant in the early and positive in the present Universe. The essence of AS quantum gravity is the presence of the Reuter fixed point whose $g$-coordinate is vanishing for large temperatures, so thus it results in a negative value for the cosmological constant if the the direction of the spiraling RG trajectories around the Reuter fixed point is the same as that of the simplest EH truncated model, a requirement fulfilled by the majority of AS quantum gravity models.
Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged. The creation of this scientific communication was supported by the University of Miskolc with funding granted to the author I. Nándori within the framework of the institution’s Scientific Excellence Support Program. (Project identifier: ME-TKTP-2025-3).
Appendix A Functional RG study of AS gravity at $T=0$
In this Appendix, we provide some details of the $T=0$ temperature functional Renormalization Group (RG) flow equations and threshold functions derived for three AS gravity models.
In order to discuss the functional RG study of CREH gravity let us rewrite the Wetterich equation using the background field method. In the background field approach the conformal factor of the background metric $\bar{g}_{\mu\nu}$ sets the physical scale of $k$. This setting is analogous to the full gravitational RG, with the exception that in this case only the conformal factor’s quantum fluctuations contribute to the running of the couplings. The path integral is taken with respect to the $\chi(x)$ quantum conformal factor field, which can be written as the sum of $\chi_{B}(x)$ fixed background field and $f(x)$ fluctuation, i.e. $\chi=\chi_{B}+f$. The effective action is functionally dependent on $\bar{f}\equiv\braket{f}$ and $\chi_{B}$, i.e. $\Gamma[\bar{f};\chi_{B}]$, or – using that the conformal factor can be given as $\phi\equiv\chi_{B}+\bar{f}$ – one can write $\Gamma[\phi,\chi_{B}]$. With the background field method the Wetterich RG equation takes the form
$k\partial_{k}\Gamma_{k}[\bar{f};\chi_{B}]=\frac{1}{2}\,\text{Tr}\,\frac{k\partial_{k}R_{k}[\chi_{B}]}{\Gamma_{k}^{(2)}[\bar{f};\chi_{B}]+R_{k}[\chi_{B}]}\,$, (13)
where $R_{k}[\chi_{B}]$ is the regulator and $\Gamma_{k}^{(2)}[\bar{f};\chi_{B}]$ is the second functional derivative with respect to $\bar{f}$ at fixed $\chi_{B}$. After one inserts an ansatz for the effective action $\Gamma_{k}[\bar{f};\chi_{B}]=\int d^{4}x\,\sqrt{\widehat{g}}{\cal L}^{\rm CREH}_{k}$ with
${\cal L}^{\rm CREH}_{k}=-\frac{3}{4\pi G_{k}}\left(-\frac{1}{2}\phi\widehat{\Box}\phi+\frac{1}{12}\widehat{\mathcal{R}}\phi^{2}-\frac{1}{6}\Lambda_{k}\phi^{4}\right),$ (14)
into the flow equation and evaluates the trace using the derivative expansion, then compares the terms on both sides the beta-functions for the couplings ($G_{k}$ and $\Lambda_{k}$) can be derived.
To compute the beta-functions for CREH gravity the following threshold functions are needed
$\Phi_{n}^{p}(w)=$ $\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\,\frac{R(y)-yR^{\prime}(y)}{[y+R(y)+w]^{p}}\,$, (15)
$\tilde{\Phi}_{n}^{p}(w)=$ $\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\,\frac{R(y)}{[y+R(y)+w]^{p}}\,.$
The anomalous dimension is
$\eta_{N}(k)=\frac{g_{k}B_{1}(\lambda_{k})}{1-g_{k}B_{2}(\lambda_{k})}\,$, (16)
where $B_{1}(\lambda_{k})$ and $B_{2}(\lambda_{k})$ functions take the form
$B_{1}(\lambda_{k})=$ $\,\frac{1}{6\pi}\,[\,\Phi_{1}^{1}(-2\lambda_{k})-\Phi_{2}^{2}(-2\lambda_{k})\,]\,$, (17)
$B_{2}(\lambda_{k})=$ $-\frac{1}{12\pi}\,[\,\tilde{\Phi}_{1}^{1}(-2\lambda_{k})-\tilde{\Phi}_{2}^{2}(-2\lambda_{k})\,]\,.$
In order to discuss the functional RG study of QEG with matter let us compute the anomalous dimension for Einstein gravity coupled to $N$-component scalar field. To do this, the functions listed below are required
$B_{1}(\lambda_{k})=$ $\,\frac{1}{6\pi}\,[\,10\,\Phi_{1}^{1}(-2\lambda_{k})+(N-8)\,\Phi_{1}^{1}(0)\,$- (18)
 $-\,36\,\Phi_{2}^{2}(-2\lambda_{k})-(12+6\,\xi\,N)\,\Phi_{2}^{2}(0)\,]\,$,
$B_{2}(\lambda_{k})=$ $\,\frac{1}{6\pi}\,[\,18\,\tilde{\Phi}_{2}^{2}(-2\lambda_{k})-5\tilde{\Phi}_{1}^{1}(-2\lambda_{k})\,]\,.$
Finally, the generalized threshold functions needed for Ghost-improved EH gravity are
$\Phi_{n}^{p,q}(w)$ $=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\frac{R(y)-yR^{\prime}(y)}{(y+R(y)+w)^{p}(y+R(y))^{q}},$
$\tilde{\Phi}_{n}^{p,q}(w)$ $=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\frac{R(y)}{(y+R(y)+w)^{p}(y+R(y))^{q}},$
$\check{\Phi}_{n}^{p,q}(w)$ $=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\frac{R^{\prime}(y)(R(y)-yR^{\prime}(y))}{(y+R(y)+w)^{p}(y+R(y))^{q}},$
$\hat{\Phi}_{n}^{p,q}(w)$ $=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dy\,y^{n-1}\frac{R(y)R^{\prime}(y)}{(y+R(y)+w)^{p}(y+R(y))^{q}}.$ (19)
One can notice that in the case of $\Phi_{n}^{p,q}(w)$ and $\tilde{\Phi}_{n}^{p,q}(w)$ the choice $q=0$ recovers the functions seen in Eqs. (15).
The anomalous dimensions of the Newton constant and the ghost wave-function renormalization are given as
$\eta_{N}$ $=\frac{gB_{1}+g^{2}\left(C_{3}C_{4}-B_{1}C_{2}\right)}{1-g\left(B_{2}+C_{2}\right)+g^{2}\left(B_{2}C_{2}-C_{1}C_{3}\right)},$ (20)
$\eta_{c}$ $=\frac{gC_{4}+g^{2}\left(B_{1}C_{1}-B_{2}C_{4}\right)}{1-g\left(B_{2}+C_{2}\right)+g^{2}\left(B_{2}C_{2}-C_{1}C_{3}\right)},$ (21)
with the functions
$B_{1}(\lambda)$ $=\frac{1}{3}(4\pi)^{1-\frac{d}{2}}\Big[d(d+1)\,\Phi^{1,0}_{\frac{d}{2}-1}-6d(d-1)\,\Phi^{2,0}_{\frac{d}{2}}$
$-4d\,\Phi^{0,1}_{\frac{d}{2}-1}-24\,\Phi^{0,2}_{\frac{d}{2}}\Big],$
$B_{2}(\lambda)$ $=-\frac{1}{6}(4\pi)^{1-\frac{d}{2}}\left[d(d+1)\,\tilde{\Phi}^{1,0}_{\frac{d}{2}-1}+6d(d-1)\,\tilde{\Phi}^{2,0}_{\frac{d}{2}}\right],$
$C_{1}(\lambda)$ $=(4\pi)^{1-\frac{d}{2}}\left[2C_{\text{gr}}\,\tilde{\Phi}^{2,1}_{\frac{d}{2}+1}-4d\left(\tilde{\Phi}^{2,2}_{\frac{d}{2}+2}+\hat{\Phi}^{2,2}_{\frac{d}{2}+2}\right)\right],$
$C_{2}(\lambda)$ $=(4\pi)^{1-\frac{d}{2}}\left[2C_{\text{gh}}\,\tilde{\Phi}^{1,2}_{\frac{d}{2}+1}+4d\left(\tilde{\Phi}^{2,2}_{\frac{d}{2}+2}+\hat{\Phi}^{2,2}_{\frac{d}{2}+2}\right)\right],$
$C_{3}(\lambda)$ $=\frac{1}{3}(4\pi)^{1-\frac{d}{2}}\left[2d\,\tilde{\Phi}^{0,1}_{\frac{d}{2}-1}+12\,\tilde{\Phi}^{0,2}_{\frac{d}{2}}\right],$
$C_{4}(\lambda)$ $=-(4\pi)^{1-\frac{d}{2}}\left[4C_{\text{gr}}\,\Phi^{2,1}_{\frac{d}{2}+1}+4C_{\text{gh}}\,\Phi^{1,2}_{\frac{d}{2}+1}\right].$ (22)
The coefficients $C_{\text{gr}}$ and $C_{\text{gh}}$ take the form
$C_{\text{gr}}=\frac{4d^{2}-9d-2}{d-2},\hskip 14.22636ptC_{\text{gh}}=\frac{2d^{2}-5d-2}{d-2}\,.$ (23)
One can recover the Einstein-Hilbert truncated anomalous dimension without ghost-improvements by setting $C_{i}(\lambda)=0$.
Appendix B Arnowitt-Deser-Misner formalism
We summarise here the main ideas behind the Arnowitt-Deser-Misner (ADM) formalism used in the discussion on the connection between spacetime foliation and thermal RG approach. In general relativity the metric is equipped with Lorentzian signature, however $\Gamma_{k}$ is given by an Euclidean path integral therefore the Lorentzian spacetime has to be recovered by Wick-rotation. QFT calculations are usually done on a fixed Minkowski background metric, which provides a clear notion of time. However, when discussing dynamical spacetimes, the role of time can be questioned. To address this the ADM-formalism is used, in which the spacetime metric $g_{\mu\nu}$ is decomposed into a lapse function $N_{l}$, a shift-vector $N_{i}$ and a metric on spatial slices $\sigma_{ij}$. These are needed, because the $4D$ spacetime is sliced into $3D$ spatial hypersurfaces with metric $\sigma_{ij}$, each hypersurface labeled with time parameter $t$. $N_{l}$ – being related to the separation between hypersurfaces – and $N_{i}$ – which is a displacement related to a point passing to the next surface – describe how to weld the hypersurfaces together to form the foliation, therefore imprinting the Euclidean spacetime with a distinguished direction. The resulting preferred time direction enables the computation of transition amplitudes from an initial to a final slice.

### v2

@@
-From negative to positive $\Lambda$ through cosmological decreasing temperatures and its connection to spacetime foliation and string theory
+From negative to positive cosmological constant through decreasing temperature of the Universe: connection with string theory and spacetime foliation results
 Abstract
-String theories naturally predict a negative, while observations on the exponential expansion of the present Universe requires a positive value for the cosmological constant $\Lambda$.
-Solution to resolve this discrepancy is known in the framework of string theory and given for Asymptotically Safe (AS) quantum gravity by using Renormalization Group (RG).
-However, the former might describes unstable worlds and the latter lacks the inclusion of temperature which is otherwise crucial in the early Universe.
-Here we present a proposal for resolving this conflict by using a modified thermal RG method where the temperature parameter $T$ is given by the inverse radius of the compactified time-like dimension, similarly to spacetime foliation.
-In our scenario not the dimensionful temperature $T$, but the dimensionless temperature $\tau=T/k$ is kept constant when the RG scale $k$ is sent to zero and string theory is assumed to take place at very high while AS quantum gravity at intermediate and low temperatures.
-We show that the modified thermal RG study of AS quantum gravity models at very high temperatures results in a negative cosmological constant while turns it into a positive parameter for low temperatures.
+[…]
 1 Introduction
@@
 This motivated attempts to resolve the discrepancy in the framework of string theory [1] by wrapping antibranes, but to get the required small and positive cosmological constant one has to wrap many fluxes and it turned out that these solutions might describe unstable worlds [2].
+[…]
+Additional attempts achieving the anti-de Sitter (AdS) – de Sitter (dS) transition include considering two interacting dark energy fluids [5], taking running Barrow entropy into account [6], or examining quintessence fields with a negative cosmological constant [7].
 Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a functional Renormalization Group (RG) method [8].
@@
 This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [17] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme.
-Moreover, in the recent work [11] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ aproach.
+Moreover, in the recent work [16] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach.
 The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [18] (i.e., non-Gaussian UV) fixed point.
@@
 Motivated by the results in [16], here we study the generality of these findings.
-We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [14, 15]) has the cosmological constant with a negative (positive, repsectively) sign for large (small) temperature.
+We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [19, 20]) has the cosmological constant with a negative (positive, respectively) sign for large (small) temperature.
 Therefore, the scheme represented in Fig.
@@
 (Project identifier: ME-TKTP-2025-3).
+This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-376) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund.
 Appendix A Functional RG study of AS gravity at $T=0$

### v3

@@
 Additional attempts achieving the anti-de Sitter (AdS) – de Sitter (dS) transition include considering two interacting dark energy fluids [6], taking running Barrow entropy into account [7], or examining quintessence fields with a negative cosmological constant [8].
-Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a functional Renormalization Group (RG) method [8].
+Another example, for the AdS–dS transition is given in [9] where the generalization of the Hawking-Page phase transition [10] is used.
+The hypothetical AdS–dS transition – which is said to alleviate the $H_{0}$ tension [11, 12, 13] and the $S_{8}$ discrepancy [14, 15, 16] – is strongly influenced by the AdS Swampland conjecture [17, 18, 19] because the logarithmic dependence ($\log|\Lambda|$) makes it impossible to cross the $\Lambda=0$ barrier at zero temperature [20, 21].
+However, it was argued and shown in [20] that at finite temperature the number of light particles can change making the sign switch possible.
+[…]
+Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a Renormalization Group (RG) in particular, the Functional Renormalization Group (FRG) method [28].
 The concept was to assume string theory at the ultraviolet (UV) and Asymptotically Safe (AS) quantum gravity & matter at the infrared (IR) scaling regimes connected by the RG scale $k$ which serves as a bridge between the negative UV and the positive IR values of the cosmological constant.
@@
 Thus, here we suggest to use the temperature instead of the RG scale $k$ to connect string theory at very high temperature and AS quantum gravity at intermediate and low temperatures of the expanding Universe.
-To achieve this one has to extend the zero-temperature functional RG method [10] to finite temperature.
+To achieve this one has to extend the zero-temperature FRG method [30] to finite temperature.
 This can be done by using the standard finite temperature extension of QFT where the time-like dimension is compactified [31, 32, 33] and its inverse radius plays the role of the dimensionful temperature parameter $T$ (in natural units).
-However, the usual choice of the perturbative RG approach, when the temperature is linked to the running (perturbative) RG scale $\mu$, i.e., $T=\mu/(2\pi)$, cannot be used in the functional RG method because one has to take the limit $k\to 0$ to obtain the quantized theory.
-Therefore, in [14] the temperature $T$ is kept constant over the RG flow, i.e., it is linked to a fixed (not running) momentum scale.
-Most of the further literature on thermal functional RG is based on this assumption.
-In this fixed $T$ approach, the dimensionful temperature is well defined and the limit $k\to 0$ can be done safely, but there is a price to pay: the RG flow equations have no fixed point solutions.
-Therefore, progress in this direction has been hampered by the fact that the usual extension of thermal functional RG method at finite temperature does not reproduce non-trivial fixed points needed for the RG analysis of various questions formulated in the framework of thermal QFT.
+In the usual finite-temperature perturbative RG approach the temperature parameter $T$ and the running (perturbative) RG scale $\mu$ are linked to each other $\mu=2\pi T$.
+The RG scales of the perturbative and the non-perturbative approaches are connected $\mu\sim k$.
+However, the $k$-dependence of the effective action of the non-perturbative FRG approach is introduced artificially to implement the Wilsonian integration of fluctuating modes and the quantised theory is obtained in the physical limit $k\to 0$ which has to be taken unavoidably.
+Thus, in the usual finite temperature FRG approach, the dimensionful temperature parameter $T$ is linked to a fixed (not running) momentum scale $\Lambda$ which results in $T=\tau\Lambda$, see [34].
+Most of the further literature on thermal FRG is based on this assumption.
+In this fixed $T$ approach, the dimensionful temperature is well defined and the limit $k\to 0$ can be done safely, but there is a price to pay: the RG flow equations have no non-trivial fixed point solutions.
 To overcome these difficulties, in Refs.
-[15, 16, 17] we proposed a modification of the usual functional RG approach af finite temperature by relating the temperature parameter to the running RG scale as $T\equiv k_{T}=\tau k$ (in natural units).
+[35, 36, 37] a modification of the usual FRG approach at finite temperature was proposed by relating the temperature parameter to the running RG scale as $T=\tau k$ (in natural units).
 In our approach
-$\tau=T/k$
-is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit.
-Thus the dimensionful parameter $T$ changes by the RG scale, so it is not considered as the temperature, but rather understood as a running cutoff for thermal fluctuation.
-The dimensionless $\tau$ is identified as the physical temperature of the surrounding plasma.
-In this case the RG flow equations have real (and not pseudo) fixed points, which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations.
+$\tau=T/k$ (1)
+is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale.
+Our working hypothesis is then to consider the dimensionful temperature parameter $T$ as a running cutoff for thermal fluctuation and to fix the parameter $\tau$ over the RG flow according (1): $T=\tau k$, where $T$ no longer plays the role of the temperature.
+We will use the notation $T\equiv k_{T}$ where $k_{T}$ is understood as the running cutoff for thermal and $k$ is for quantum fluctuations.
+In our approach the “true” temperature (the one entering the thermodynamic quantities such as the free energy that one aims at calculating) is proportional to the dimensionless quantity $\tau$ which is kept constant over the RG flow; it can be chosen arbitrarily, it is not restricted to a specific value and therefore can have any finite value.
+The relation between $\tau$ and the “true” temperature $T_{true}$ is $\tau=k_{B}T_{true}/J$, where $J$ is an energy scale of the system depending on the microscopic details of the system.
+Since our results depend only on $\tau$ and not separately on $T_{true}$ and $J$, we keep the discussion in terms of $\tau$.
+In other words, one can say $\tau$ measures how large the thermal fluctuations are compared to the quantum ones.
+Thus, our thermal RG scheme differs from the usual non-perturbative and also from the perturbative ones.
+The rest of the paper is devoted to understand the consequences of our working hypothesis on the thermal evolution of the cosmological constant.
 In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points.
@@
 This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [37] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme.
-Moreover, in the recent work [16] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach.
+Moreover, in the recent work [36] we studied the thermal RG of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach.
 The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [38] (i.e., non-Gaussian UV) fixed point.
@@
 Figure 1: Schematic figure of the time evolution (temperature, i.e., $\tau$-dependence) of the scale factor $R(t)$ and the cosmological constant $\Lambda(t)$.
-Motivated by the results in [16], here we study the generality of these findings.
+Motivated by the results of [36], here we study the generality of these findings.
 We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [39, 40]) has the cosmological constant with a negative (positive, respectively) sign for large (small) temperature.
@@
 The harmonic gauge choice is also used in order to compare the result to the EH truncated gravity without ghost-improvement.
-Beta-functions in $d$ dimensions are found to be
+Beta-functions in $d$ dimensions are
 $\beta_{g}=$ $\,(d-2+\eta_{N})\,g_{k}\,$, (10)
@@
 In all three gravitational models we found that the $g$-component of the Reuter fixed point ($g^{*}$) disappears as the temperature increases, i.e., with $\tau\to\infty$, see Fig. 2.
-Figure 2: Thermal RG flow diagrams of CREH, QEG with $N$-component scalar field (with chosen parameters $\xi=10$ and $N=2$) and Ghost-improved EH gravity for various $\tau$ dimensionless temperature values.
+Figure 2: Thermal RG flow diagrams of CREH, QEG with $N$-component scalar field (with $\xi=10$ and $N=2$) and Ghost-improved EH gravity for various $\tau$ dimensionless temperature values.
 With $\tau\to\infty$ the $g^{*}\to 0$ limit is reached in each case.
@@
 The essence of AS quantum gravity is the presence of the Reuter fixed point whose $g$-coordinate is vanishing for large temperatures, so thus it results in a negative value for the cosmological constant if the the direction of the spiraling RG trajectories around the Reuter fixed point is the same as that of the simplest EH truncated model, a requirement fulfilled by the majority of AS quantum gravity models.
+[…]
 Acknowledgements
 The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged.
-The creation of this scientific communication was supported by the University of Miskolc with funding granted to the author I.
-Nándori within the framework of the institution’s Scientific Excellence Support Program.
-(Project identifier: ME-TKTP-2025-3).
 This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-376) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund.
+Note added
+After the submission several interesting papers appeared [26, 27, 57, 58, 59, 60, 61], where the authors discuss observational results in terms of a phenomenological model supporting a scenario compatible with our results.
+Further investigation on this point is currently on-going.
 Appendix A Functional RG study of AS gravity at $T=0$
@@
 $C_{\text{gr}}=\frac{4d^{2}-9d-2}{d-2},\hskip 14.22636ptC_{\text{gh}}=\frac{2d^{2}-5d-2}{d-2}\,.$ (24)
-One can recover the Einstein-Hilbert truncated anomalous dimension without ghost-improvements by setting $C_{i}(\lambda)=0$.
+One can recover the Einstein-Hilbert truncated anomalous dimension without ghost-improvements by setting $C_{i}=0$.
 Appendix B Arnowitt-Deser-Misner formalism
-We summarise here the main ideas behind the Arnowitt-Deser-Misner (ADM) formalism used in the discussion on the connection between spacetime foliation and thermal RG approach.
-In general relativity the metric is equipped with Lorentzian signature, however $\Gamma_{k}$ is given by an Euclidean path integral therefore the Lorentzian spacetime has to be recovered by Wick-rotation.
-QFT calculations are usually done on a fixed Minkowski background metric, which provides a clear notion of time.
+The Arnowitt-Deser-Misner (ADM) formalism is used to connect spacetime foliation and thermal RG.
+In general relativity the metric is equipped with Lorentzian signature, however $\Gamma_{k}$ is given by an Euclidean path integral thus the Lorentzian spacetime has to be recovered by Wick-rotation.
+QFT calculations are usually done on a fixed Minkowski background which provides a clear notion of time.
 However, when discussing dynamical spacetimes, the role of time can be questioned.
@@
 These are needed, because the $4D$ spacetime is sliced into $3D$ spatial hypersurfaces with metric $\sigma_{ij}$, each hypersurface labeled with time parameter $t$.
-$N_{l}$ – being related to the separation between hypersurfaces – and $N_{i}$ – which is a displacement related to a point passing to the next surface – describe how to weld the hypersurfaces together to form the foliation, therefore imprinting the Euclidean spacetime with a distinguished direction.
+$N_{l}$ – being related to the separation between hypersurfaces – and $N_{i}$ – which is a displacement related to a point passing to the next surface – describe how to weld the hypersurfaces together to form the foliation, thus imprinting the Euclidean spacetime with a distinguished direction.
 The resulting preferred time direction enables the computation of transition amplitudes from an initial to a final slice.